# Privacy-Aware Sequential Learning

### v4

@@ 1 Introduction
To address these questions, we adopt the metric differential privacy (mDP) framework (Chatzikokolakis et al. 2013) to study the behavior of privacy-aware agents. In our setting, agents endogenously add noise to their actions to limit inferences from other observers, thereby altering the dynamics of learning. Contrary to the common intuition that additional noise should reduce the informativeness of public observations (Le et al. 2017, Papachristou and Rahimian 2025), we show that appropriately designed, privacy-motivated randomization can enhance learning by asymmetrically amplifying the log-likelihood ratio across states, thus increasing the information conveyed by each action. This asymmetry stems from a strategic tradeoff between privacy and utility: agents randomize more when their actions would otherwise reveal sensitive information, and less when the risk of revelation is low. Unlike prior work that ties learning speed solely to the structure of private signals (Hann-Caruthers et al. 2018, Rosenberg and Vieille 2019), we demonstrate that privacy-driven noise reshapes aggregation dynamics and can accelerate convergence.
@@ 1.1 Main related work
Our study contributes to the broader literature on how privacy constraints affect learning and decision-making. In the context of sequential learning, Tsitsiklis et al. (2021) and Xu (2018) analyze the query complexity of private learning and show that stronger privacy constraints require more information acquisition for accurate inference. Although their models share structural similarities with ours, they focus on independent queries, whereas we study path-dependent decision-making in which each agent’s action shapes the information available to others. In distributed settings, privacy-preserving noise is typically modeled as exogenous (Papachristou and Rahimian 2025, Rizk et al. 2023, Tao et al. 2023), whereas we treat it as an endogenous choice. We further allow agents to differ in their privacy budgets, consistent with evidence of heterogeneous privacy preferences (Acquisti and Grossklags 2005) and with personalized/heterogeneous DP frameworks (Alaggan et al. 2015, Acharya et al. 2025) This shift from algorithmic to behavioral modeling highlights how rational agents strategically balance privacy and learning performance, offering new insights at the intersection of privacy theory and social learning. To our knowledge, the only work examining social learning in privacy-preserving data collection is Akbay et al. (2021), which focuses on simultaneous actions and does not address sequential learning or information cascades.
@@ 1.2 Main contribution
Probability of correct cascade $\frac{p^{k}-1}{p^{2k}-1}$ $\frac{\rho(\varepsilon)^{k-1}}{\rho(\varepsilon)^{2k-1}}$ $\frac{\bar{\rho}^{k-1}}{\bar{\rho}^{2k-1}}$
Information cascade threshold $k=2$ $k=\left\lfloor\log_{\frac{\rho(\varepsilon)}{1-\rho(\varepsilon)}}\frac{p}{1-p}\right\rfloor+1$ $k=\left\lfloor\log_{\frac{\bar{\rho}}{1-\bar{\rho}}}\frac{p}{1-p}\right\rfloor+1$
@@
Table 1: Summary of theoretical results under different privacy settings. Here, $\rho(\varepsilon)=u(\varepsilon)(1-p)+p(1-u(\varepsilon))$, $\bar{\rho}=\bar{u}(1-p)+p(1-\bar{u})$, $\bar{u}=E_{\varepsilon}[u(\varepsilon)]=E_{\varepsilon}\left[\frac{1}{1+e^{\varepsilon}}\right]$, $C(\varepsilon)=\frac{\varepsilon\sigma^{2}}{4}\left(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}\right)$, and $C_{1}$ and $\tilde{C}$ are constants. The total number of incorrect actions differs from the time of the first correct action by a constant and is omitted.
@@
The remainder of this paper is organized as follows. First, we formally define our problem setup in Section 2, including the sequential decision making framework and integration of privacy constraints. Section 3.1 focuses on the case of binary signals, where we analyze how differential privacy affects the probability of correct cascades using a Markov chain representation. In Section 4.1, we extend our analysis to continuous signals, introducing the smooth randomized response strategy and showing how it accelerates and improves learning efficiency. Finally, Section 5 summarizes our findings and discusses potential future research directions, including broader implications for privacy-aware decision making in sequential learning environments.
@@ 2 A Model of Sequential Learning with Privacy
We consider a classical sequential learning model: There exists an unknown binary state of nature, denoted by $\theta\in\{-1,+1\}$, with the same prior probability $1/2$; each agent, indexed by $n\in\{1,2,\ldots\}$, receives a private signal $s_{n}\in\mathscr{X}$, drawn independently and identically distributed (i.i.d.) from a conditional distribution $F_{\theta}$, that is, $s_{n}\sim F_{\theta}$, where $F_{\theta}$ depends on the true state $\theta$. Without considering privacy concerns, each agent $n$ takes an action $a_{n}\in\{-1,+1\}$ based on the information received, including the history of previous actions denoted by $h_{n-1}:=\{a_{1},\cdots,a_{n-1}\}$ and their own private signal $s_{n}$. If the action $a_{n}$ matches $\theta$, the agent’s utility is 1; otherwise, it is 0.
@@
Based on their private signal and the observed history of prior adoption decisions, each agent chooses a true adoption decision $a_{n}$. To protect privacy, the agent may then misreport via a randomized mechanism $\mathcal{M}$ that injects bounded noise into the publicly observed action. This reporting is not fully random: agents internalize the value of informative records—for example, for eligibility verification, booster scheduling, adverse-event monitoring, and care coordination—and they may face moral, reputational, or regulatory costs from blatant misreporting. Accordingly, $\mathcal{M}$ is calibrated by a fixed privacy budget to balance privacy concerns with public informativeness. Formally, agent $n$ selects a reporting strategy to maximize their expected utility, subject to a fixed privacy budget:
@@ 3.1 Main Results for Binary Signals
Theorem 3.4 (Randomized Response Strategy for Binary Model)
@@
To illustrate, consider the vaccine adoption scenario discussed earlier, in which individuals decide sequentially whether to take a new vaccine, and each person observes aggregate adoption statistics before making their decision. In the early stages, individuals rely on their private health-related signals and can apply randomized response techniques to protect sensitive medical information, such as reporting adoption with probability $p$ even if they privately choose not to adopt and vice versa. This strategy ensures plausible deniability and guards against inference of personal health conditions, especially when the vaccine’s risks or effectiveness are controversial.
@@
Consider a correct signal probability $p$ and an information cascade threshold $k$, and let $\alpha=(1-p)/p$, $v_{k}=\frac{1-\alpha^{\frac{k-2}{k-1}}}{1-\alpha^{\frac{k-2}{k-1}}+\alpha^{\frac{-1}{k-1}}-\alpha}$, $\varepsilon_{k}=\log\frac{1-v_{k}}{v_{k}}$ and $k=\lfloor\log_{\frac{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}}\frac{p}{(1-p)}\rfloor+1$. Given the privacy budget $\varepsilon$, agents employ the randomized response strategy with flip probability $u_{n}=u(\varepsilon)=\frac{1}{1+e^{\varepsilon}}$, then for $\varepsilon\in(\varepsilon_{k+1},\varepsilon_{k}]$, the information cascade threshold $k$ does not change, and the probability of the correct cascade increases with $\varepsilon$.
@@
where $\rho(\varepsilon)=\frac{1-\tilde{u}}{\tilde{u}}$ and $\tilde{u}=u(1-p)+p(1-u)$. The associated cascade threshold is given in Lemma 8.1:
@@
Consider a correct signal probability $p$ and an information cascade threshold $k$, and define $\alpha=(1-p)/p$. Let the threshold-related flip probability be $\tilde{v}_{k}=\frac{1-\alpha^{\frac{k-2}{k-1}}}{1-\alpha^{\frac{k-2}{k-1}}+\alpha^{\frac{-1}{k-1}}-\alpha}$, and let the cascade threshold satisfy $k=\left\lfloor\log_{\frac{\bar{u}(1-p)+p(1-\bar{u})}{(1-\bar{u})(1-p)+\bar{u}p}}\frac{p}{1-p}\right\rfloor+1$. Given a privacy budget $\varepsilon_{n}$ for each agent, agents adopt the randomized response strategy with flipping probability $u_{n}(\varepsilon_{n})=\frac{1}{1+e^{\varepsilon_{n}}}$, then for $\bar{u}\in[\tilde{v}_{k},\tilde{v}_{k+1})$, the cascade threshold $k$ remains unchanged, and the probability of a correct cascade decreases with expected flipping probability $\bar{u}$.
@@ 4 Learning Efficiency with Continuous Signals
As illustrated in Figure 2, for each value of the public belief $\ell$, there exists a unique signal threshold $t(\ell)$ such that agent $n$’s optimal action switches from 1-1 to +1+1 as their private signal $s_{n}$ crosses this threshold. This indicates that decisions are most sensitive to noise near this threshold. However, the $\varepsilon$-DP requirement enforces indistinguishability across all signal pairs, which inevitably includes pairs on opposite sides of the decision boundary. As a result, agents must randomize their actions with the same fixed probability, regardless of the signal’s informativeness. We formalize this observation as follows:
Theorem 4.2 (Randomized Response Strategy for Continuous Model)
In the sequential learning model with Gaussian signals under an $\varepsilon$-differential privacy constraint, agents adopt a randomized response strategy with a fixed flip probability $u_{n}=u=\frac{1}{1+e^{\varepsilon}}$, independent of the agent index $n$.
Theorem 4.3 (Failure of Asymptotic Learning)
@@ 4.1 Main Results for Continuous Signals
Theorem 4.10 (Smoothly Asymptotic Learning)
@@
The $\Theta_{\varepsilon}(\log(n))$ asymptotic learning rate in Theorem 4.11 represents an improvement by an order of $\sqrt{\log(n)}$ over the $\Theta(\sqrt{\log(n)})$ asymptotic learning speed established by Hann-Caruthers et al. (2018) in the non-private regime. To better understand this improvement, assume $\theta=+1$. For Gaussian signals, asymptotic learning is guaranteed, which implies that over time, more agents will take action +1+1 than 1-1. To protect their private signals, agents use the smooth randomized response to add noise to their actions. Importantly, this noise is asymmetric: In the true state, where signals are more aligned with the majority, agents require less distortion to maintain privacy; under the false state, where signals are more likely to contradict the majority, agents need to introduce more noise. This asymmetry arises because privacy protection requires obscuring how surprising a signal is — more surprising signals (those against the majority) require more distortion. As a result, actions under the correct state remain more stable, while those under the wrong state are noisier. This amplifies the difference between the two states in terms of observed behavior and makes actions more informative for later agents, thereby accelerating learning.
@@
where $C_{1}$ is a positive constant that does not depend on $\varepsilon$. The series in Equation 7 converges if, and only if, $\frac{2}{\varepsilon\sigma^{2}}>1$; and thus, the expected time to the first correct action is finite $(\mathbb{E}[\tau]<+\infty)$, if, and only if, $\frac{2}{\varepsilon\sigma^{2}}>1$.
@@
A martingale argument shows that $\mathbb{P}(x_{k}=+1|l_{k},\theta=+1)$ is positive; therefore, we have $u_{n}\sim 1-\pi_{n+1}^{*}$. Since $l_{n}=\log\left(\frac{\pi_{n}^{*}}{1-\pi_{n}^{*}}\right)$, we have $1-\pi_{n}^{*}\sim e^{-l_{n}}$, where $l_{n}=\frac{2}{\varepsilon\sigma^{2}}\ln(C(\varepsilon)n)+c$ where $c$ is a constant. Thus, the expected time $\mathbb{E}[\tau]$ until the first correct action satisfies
@@
for some constant $C(\varepsilon)$ depending on $\varepsilon$. Since $\frac{2}{\varepsilon\sigma^{2}}>1$, the series converges and thus the expected time until the first correct action is finite.
@@
where $C_{2}$ is a positive constant that does not depend on $\varepsilon$. The series in Equation 11 converges if, and only if, $\frac{2}{\varepsilon\sigma^{2}}>1$; and thus, the expected total number of incorrect actions is finite, $\mathbb{E}[W]<+\infty$, if, and only if, $\frac{2}{\varepsilon\sigma^{2}}>1$.
@@
$\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}|\mathcal{H}_{\tau_{k-1}}]\leq\frac{\tau_{k-1}}{1-\tau_{k-1}}\mathbb{E}[\tau]\leq\mathbb{E}[\tau]$
@@
When $\frac{2}{\varepsilon\sigma^{2}}>1$, both the expected time to the first correct action and the expected total number of incorrect actions are finite. This raises a natural question: what is the optimal privacy budget $\varepsilon^{*}$ that minimizes these two expectations? Both quantities are determined by the expression $C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}}$, and our goal is to find the value of $\varepsilon^{*}$ that minimizes this term. However, due to the involvement of the Riemann zeta function in the infinite sum, obtaining a closed-form solution is infeasible. Therefore, in Figure 4 we fine the optimal privacy budget $\varepsilon^{*}\approx 0.76$ numerically. While $\varepsilon$ is endogenously chosen by agents in equilibrium, this calculation serves as a normative benchmark for a planner or platform that can influence privacy choices through incentives or design. Such a benchmark helps quantify the potential welfare loss from purely self-interested privacy decisions and highlights the scope for policy interventions that align individual incentives with socially optimal information aggregation, echoing broader discussions on principled $\varepsilon$-selection in both practice (Dwork et al. 2019) and economics (Hsu et al. 2014).
Figure 4: The value of $C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}}$ as a function of the privacy budget $\varepsilon$ ($\sigma=1$).
@@
In the homogeneous setting, we have already characterized the optimal privacy budget $\varepsilon^{*}$ for minimizing both the expected time to the first correct action and the expected total number of incorrect actions, assuming a fixed and positive privacy budget $\varepsilon$. However, the optimal convergence rate of learning in this setting remains unknown when $\varepsilon$ approaches zero. Interestingly, as $\varepsilon$ approaches zero, the speed of learning can increase. To investigate this phenomenon, we consider a heterogeneous setting in which each agent’s privacy budget $\varepsilon_{n}$ is drawn from a distribution that assigns positive probability mass to values of $\varepsilon$ close to zero. This captures agents with varying privacy concerns. We assume $\varepsilon_{n}\sim U[0,1]$ for simplicity, noting that the results extend to $\varepsilon_{n}\sim U[0,a]$ for any $a>0$ after appropriate rescaling. Additional simulation results for different choices of privacy budget intervals and their corresponding learning performance are provided in Appendix LABEL:app:simulation_heterogeneous. We show that the convergence rate of the public log-likelihood ratio can increase to $\Theta(\sqrt{n})$ due to the presence of agents with strong privacy preferences. Before presenting the formal results, we first establish that agents with $\varepsilon_{n}\sim U[0,1]$ still achieve asymptotic learning, as shown in Theorem 4.14.
@@
Theorem 4.16 (Learning Rate Bounds under Heterogeneous Privacy Budgets)
@@
Theorem 4.16 highlights a fundamental limit in privacy-aware sequential learning: the presence of agents with very small privacy budgets effectively determines the maximal rate at which public belief can converge. To place this result in context, it is useful to compare it with two benchmark settings. In the ideal case where agents directly observe i.i.d. Gaussian signals without any sequential decision-making, the convergence rate of the aggregated belief is $\Theta(n)$, corresponding to the classical law of large numbers. In the standard non-private sequential learning setting (i.e., without any privacy-preserving behavior), the convergence rate is reduced to $\Theta(\sqrt{\log(n)})$ due to the information loss from agents observing only the actions of their predecessors (Hann-Caruthers et al. 2018, Rosenberg and Vieille 2019).
Our result shows that, when smooth randomized response is employed under heterogeneous privacy budgets, the convergence rate can improve to $\Theta(\sqrt{n})$, which is strictly faster than that in the non-private sequential setting. This somewhat surprising outcome arises from two key factors: the presence of highly privacy-conscious agents (with small $\varepsilon_{n}$) and the inherent asymmetry of the smooth randomized response strategy under the two possible states. Specifically, under the true state, the average probability of flipping the action is higher than under the false state, which makes the observed actions more informative in expectation. Nonetheless, Theorem 4.16 confirms that $\Theta(\sqrt{n})$ is the fastest achievable convergence rate under any distribution of privacy budgets, and thus represents the optimal learning rate under privacy constraints in sequential settings.

### v5

@@ 1 Introduction
To address these questions, we adopt the metric differential privacy (mDP) framework (Chatzikokolakis et al. 2013) to study the behavior of privacy-aware agents. In our setting, agents endogenously add noise to their actions to limit inferences from other observers, thereby altering the dynamics of learning. Contrary to the common intuition that additional noise should reduce the informativeness of public observations (Le et al. 2017, Papachristou and Rahimian 2025), we show that privacy-motivated randomization can enhance learning by asymmetrically amplifying the log-likelihood ratio across states, thus increasing the information conveyed by each action. This asymmetry stems from a strategic tradeoff between privacy and utility: agents randomize more when their actions would otherwise reveal sensitive information, and less when the risk of revelation is low. Unlike prior work that ties learning speed solely to the structure of private signals (Hann-Caruthers et al. 2018, Rosenberg and Vieille 2019), we demonstrate that privacy-driven noise reshapes aggregation dynamics and can accelerate convergence.
@@ 1.1 Main related work
Our study contributes to the broader literature on how privacy constraints affect learning and decision-making. In the context of sequential learning, Tsitsiklis et al. (2021) and Xu (2018) analyze the query complexity of private learning and show that stronger privacy constraints require more information acquisition for accurate inference. Although their models share structural similarities with ours, they focus on independent queries, whereas we study path-dependent decision-making in which each agent’s action shapes the information available to others. In distributed settings, privacy-preserving noise is typically modeled as exogenous (Papachristou and Rahimian 2025, Rizk et al. 2023, Tao et al. 2023), whereas we treat it as an endogenous choice. We further allow agents to differ in their privacy budgets, consistent with evidence of heterogeneous privacy preferences (Acquisti and Grossklags 2005) and with personalized/heterogeneous DP frameworks (Alaggan et al. 2015, Acharya et al. 2025). This shift from algorithmic to behavioral modeling highlights how rational agents strategically balance privacy and learning performance, offering new insights at the intersection of privacy theory and social learning. To our knowledge, the only work examining social learning in privacy-preserving data collection is Akbay et al. (2021), which focuses on simultaneous actions and does not address sequential learning or information cascades.
@@ 1.2 Main contribution
Probability of correct cascades $\frac{\rho^{k}-1}{\rho^{2k}-1}$ $\frac{\rho(\varepsilon)^{k}-1}{\rho(\varepsilon)^{2k}-1}$ $\frac{\bar{\rho}^{k}-1}{\bar{\rho}^{2k}-1}$
Information cascade threshold $k=2$ $k=\left\lfloor\log_{\rho(\varepsilon)}\frac{1-p}{p}\right\rfloor+1$ $k=\left\lfloor\log_{\bar{\rho}}\frac{1-p}{p}\right\rfloor+1$
@@
Table 1: Summary of theoretical results under different privacy settings. Here, $\rho=\frac{1-p}{p},\rho(\varepsilon)=\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}$, $\bar{\rho}=\frac{(1-\bar{u})(1-p)+\bar{u}p}{\bar{u}(1-p)+p(1-\bar{u})}$, $\bar{u}=E_{\varepsilon}[u(\varepsilon)]=E_{\varepsilon}\left[\frac{1}{1+e^{\varepsilon}}\right]$, $C(\varepsilon)=\frac{\varepsilon\sigma^{2}}{4}\left(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}\right)$, and $C_{1}$ and $\tilde{C}$ are constants. The total number of incorrect actions differs from the time of the first correct action by a constant factor and is omitted.
@@
The remainder of this paper is organized as follows. First, we formally define our problem setup in Section 2, including the sequential decision making framework and integration of privacy constraints. Section 3 focuses on the case of binary signals, where we analyze how differential privacy affects the probability of correct cascades using a Markov chain representation. In Section 4, we extend our analysis to continuous signals, introducing the smooth randomized response strategy and showing how it accelerates and improves learning efficiency. Finally, Section 5 summarizes our findings and discusses potential future research directions, including broader implications for privacy-aware decision making in sequential learning environments.
@@ 2 A Model of Sequential Learning with Privacy
We consider a classical sequential learning model: There exists an unknown binary state of nature, denoted by $\theta\in\{-1,+1\}$, with the same prior probability $1/2$; each agent, indexed by $n\in\{1,2,\ldots\}$, receives a private signal $s_{n}\in\mathscr{X}$, drawn independently and identically distributed (i.i.d.) from a conditional distribution $F_{\theta}$, that is, $s_{n}\sim F_{\theta}$, where $F_{\theta}$ depends on the true state $\theta$. Without considering privacy concerns, each agent $n$ takes an action $a_{n}\in\{-1,+1\}$ based on the information received, including the history of previous reported actions denoted by $h_{n-1}:=\{x_{1},\cdots,x_{n-1}\}$, where $x_{i}$ is the reported action of agent $i$, and their own private signal $s_{n}$.
@@
Based on their private signal and the observed history of prior adoption decisions, each agent chooses a true adoption decision $a_{n}$. To protect privacy, the agent may then misreport via a randomized mechanism $\mathcal{M}$ that injects bounded noise into the publicly observed action. This reporting is not fully random: agents internalize the value of informative records—for example, for eligibility verification, booster scheduling, adverse-event monitoring, and care coordination—and they may face moral, reputational, or regulatory costs from blatant misreporting. Thus, If the reported action $x_{n}$ matches $\theta$, the agent’s utility is 1 ; otherwise, it is 0. Accordingly, $\mathcal{M}$ is calibrated by a fixed privacy budget to balance privacy concerns with public informativeness. Formally, agent $n$ selects a reporting strategy to maximize their expected utility, subject to a fixed privacy budget:
@@ 3.1 Main Results for Binary Signals
Theorem 3.4 (Randomized Response Strategy for the Binary Model)
@@
To illustrate, consider the vaccine adoption scenario discussed earlier, in which individuals decide sequentially whether to take a new vaccine, and each person observes aggregate adoption statistics before making their decision. In the early stages, individuals rely on their private health-related signals and can apply randomized response techniques to protect sensitive medical information, such as reporting adoption with probability $u_{n}$ even if they privately choose not to adopt and vice versa. This strategy ensures plausible deniability and guards against inference of personal health conditions, especially when the vaccine’s risks or effectiveness are controversial.
@@
Consider a correct signal probability $p$ and privacy budget $\epsilon$, and define $\alpha=(1-p)/p$ and $u(\varepsilon)=\frac{1}{1+e^{\varepsilon}}$. Let $v_{k}=\frac{1-\alpha^{\frac{k-2}{k-1}}}{1-\alpha^{\frac{k-2}{k-1}}+\alpha^{\frac{-1}{k-1}}-\alpha}$ and $\varepsilon_{k}=\log\frac{1-v_{k}}{v_{k}}$, where $k$ is the information cascade threshold given by $k=\lfloor\log_{\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}}\frac{1-p}{p}\rfloor+1$. Subject to the privacy budget $\varepsilon$, agents employ the randomized response strategy with flip probability $u_{n}=u(\varepsilon)$; subsequently, for $\varepsilon\in(\varepsilon_{k+1},\varepsilon_{k}]$, the information cascade threshold $k$ does not change and the probability of the correct cascade increases with $\varepsilon$.
@@
where $\rho(\varepsilon)=\frac{(1-u(\varepsilon))(1-p)+pu(\varepsilon)}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}$. The associated cascade threshold is given in Lemma 8.1:
@@
Consider a correct signal probability $p$ and privacy budget $\varepsilon_{n}$, and define $\alpha=(1-p)/p$ and $\bar{u}=\mathbb{E}_{\varepsilon_{n}}[\frac{1}{1+e^{\varepsilon_{n}}}]$. Let $\tilde{v}_{k}=\frac{1-\alpha^{\frac{k-2}{k-1}}}{1-\alpha^{\frac{k-2}{k-1}}+\alpha^{\frac{-1}{k-1}}-\alpha}$, where $k$ is information cascade threshold given by $k=\left\lfloor\log_{\frac{(1-\bar{u})(1-p)+\bar{u}p}{\bar{u}(1-p)+p(1-\bar{u})}}\frac{1-p}{p}\right\rfloor+1$. Subject to a privacy budget $\varepsilon_{n}$ for each agent, agents adopt the randomized response strategy with flipping probability $u_{n}(\varepsilon_{n})=\frac{1}{1+e^{\varepsilon_{n}}}$, then for $\bar{u}\in[\tilde{v}_{k},\tilde{v}_{k+1})$, the cascade threshold $k$ remains unchanged, and the probability of a correct cascade decreases with expected flipping probability $\bar{u}$.
@@ 4 Learning Efficiency with Continuous Signals
As illustrated in Figure 2, for each value of the public belief $\ell$, there exists a unique signal threshold $t(\ell)$ such that agent $n$’s optimal action switches from 1-1 to +1+1 as their private signal $s_{n}$ crosses this threshold. This indicates that decisions are most sensitive to noise near this threshold. However, the $\varepsilon$-DP requirement enforces indistinguishability across all signal pairs, which inevitably includes pairs on opposite sides of the decision boundary. As a result, agents must randomize their actions with the same fixed probability, regardless of the signal’s informativeness. We provide the detailed formal analysis in Appendix 9.
@@ 4.1 Main Results for Continuous Signals
Theorem 4.8 (Asymptotic Learning with Smooth Randomized Response)
@@
The $\Theta_{\varepsilon}(\log(n))$ asymptotic learning rate in Theorem 4.9 represents an improvement by an order of $\sqrt{\log(n)}$ over the $\Theta(\sqrt{\log(n)})$ asymptotic learning speed established by Hann-Caruthers et al. (2018) in the non-private regime. To better understand this improvement, assume that $\theta=+1$. For Gaussian signals, asymptotic learning is guaranteed, which implies that over time, more agents will take action +1+1 than 1-1. To protect their private signals, agents use the smooth randomized response to add noise to their actions. Importantly, this noise is asymmetric: In the true state, where signals are more aligned with the majority, agents require less distortion to maintain privacy; under the false state, where signals are more likely to contradict the majority, agents need to introduce more noise. This asymmetry arises because privacy protection requires obscuring how surprising a signal is — more surprising signals (those against the majority) require more distortion. As a result, actions under the correct state remain more stable, while those under the wrong state are noisier. This amplifies the difference between the two states in terms of observed behavior and makes actions more informative for later agents, thereby accelerating learning.
@@
where $C_{1}$ is a positive constant that does not depend on $\varepsilon$. The series in Equation 7 converges if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$; and thus, the expected time to the first correct action is finite $(\mathbb{E}[\tau]<+\infty)$, if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$.
@@
A martingale argument shows that $\mathbb{P}(x_{k}=+1|l_{k},\theta=+1)$ is positive and $\pi_{n+1}^{*}\geq\frac{1}{2}$; therefore, we have $u_{n}\sim 1-\pi_{n+1}^{*}$. Since $l_{n}=\log\left(\frac{\pi_{n}^{*}}{1-\pi_{n}^{*}}\right)$, we have $1-\pi_{n}^{*}\sim e^{-l_{n}}$, where $l_{n}=\frac{2}{\varepsilon\sigma^{2}}\ln(C(\varepsilon)n)+c$ where $c$ is a constant. Thus, the expected time $\mathbb{E}[\tau]$ until the first correct action satisfies
@@
for some constant $C(\varepsilon)$ depending on $\varepsilon$. Since $\varepsilon<\frac{2}{\sigma^{2}}$, the series converges and thus the expected time until the first correct action is finite.
@@
where $C_{2}$ is a positive constant that does not depend on $\varepsilon$. The series in Equation 11 converges if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$; and thus, the expected total number of incorrect actions is finite, $\mathbb{E}[W]<+\infty$, if, and only if, $\varepsilon<\frac{2}{\sigma^{2}}$.
@@
$\mathbb{E}_{-}[\Delta_{k}\mathbf{1}_{\sigma_{k}<+\infty}|\mathcal{H}_{\tau_{k-1}}]\leq\frac{\pi_{\tau_{k-1}}}{1-\pi_{\tau_{k-1}}}\mathbb{E}[\tau]\leq\mathbb{E}[\tau]$
@@
When $\varepsilon<\frac{2}{\sigma^{2}}$, both the expected time to the first correct action and the expected total number of incorrect actions are finite. This raises a natural question: what is the optimal privacy budget $\varepsilon^{*}$ that minimizes these two expectations? Both quantities are determined by the expression $C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}}$, and our goal is to find the value of $\varepsilon^{*}$ that minimizes this term. However, due to the involvement of the Riemann zeta function in the infinite sum, obtaining a closed-form solution is infeasible. Therefore, we set $\sigma=\sqrt{2}$ so that $\tfrac{2}{\sigma^{2}}=1$, which normalizes the range of $\varepsilon$ to $(0,1)$. Accordingly, in Figure 4 we numerically obtain the optimal privacy budget $\varepsilon^{*}\approx 0.76$. While $\varepsilon$ is endogenously chosen by agents in equilibrium, this calculation serves as a normative benchmark for a planner or platform that can influence privacy choices through incentives or design. Such a benchmark helps quantify the potential welfare loss from purely self-interested privacy decisions and highlights the scope for policy interventions that align individual incentives with socially optimal information aggregation, echoing broader discussions on principled $\varepsilon$-selection in both practice (Dwork et al. 2019) and economics (Hsu et al. 2014).
Figure 4: The value of $C(\varepsilon)^{-\frac{2}{\varepsilon\sigma^{2}}}\sum_{n=1}^{\infty}n^{-\frac{2}{\varepsilon\sigma^{2}}}$ as a function of the privacy budget $\varepsilon$ ($\sigma=\sqrt{2}$).
@@
In the homogeneous setting, we have already characterized the optimal privacy budget $\varepsilon^{*}$ for minimizing both the expected time to the first correct action and the expected total number of incorrect actions, assuming a fixed and positive privacy budget $\varepsilon$. However, the optimal convergence rate of learning in this setting remains unknown when $\varepsilon$ approaches zero. Interestingly, as $\varepsilon$ approaches zero, the speed of learning can increase. To investigate this phenomenon, we consider a heterogeneous setting in which each agent’s privacy budget $\varepsilon_{n}$ is drawn from a distribution that assigns positive probability mass to values of $\varepsilon$ close to zero. This captures agents with varying privacy concerns. We assume $\varepsilon_{n}\sim U[0,1]$ for simplicity, noting that the results extend to $\varepsilon_{n}\sim U[0,a]$ for any $a>0$ after appropriate rescaling. Additional simulation results for different choices of privacy budget intervals and their corresponding learning performance are provided in Figure 5. We show that the convergence rate of the public log-likelihood ratio can increase to $\Theta(\sqrt{n})$ due to the presence of agents with strong privacy preferences. Before presenting the formal results, we first establish that agents with $\varepsilon_{n}\sim U[0,1]$ still achieve asymptotic learning, as shown in Theorem 4.12.
@@
Theorem 4.14 (Learning Rate Bound under Heterogeneous Privacy Budgets)
@@
Theorem 4.14 highlights a fundamental limit in privacy-aware sequential learning: the presence of agents with very small privacy budgets effectively determines the maximal rate at which public belief can converge. To place this result in context, it is useful to compare it with the following two benchmarks. In the ideal case where an agent directly observe $n$ i.i.d. Gaussian private signals, the convergence rate of the aggregated belief is $\Theta(n)$, corresponding to the classical law of large numbers. In the standard non-private sequential learning setting (i.e., without any privacy-preserving behavior), the convergence rate is reduced to $\Theta(\sqrt{\log(n)})$ due to the information loss from agents observing only the actions of their predecessors (Hann-Caruthers et al. 2018, Rosenberg and Vieille 2019).
Our result shows that, when smooth randomized response is employed under heterogeneous privacy budgets, the convergence rate can improve to $\Theta(\sqrt{n})$, which is strictly faster than that in the non-private sequential setting. This somewhat surprising outcome arises from two key factors: the presence of highly privacy-conscious agents (with small $\varepsilon_{n}$) and the inherent asymmetry of the smooth randomized response strategy under the two possible states. Specifically, under the true state, the average probability of flipping the action is lower than under the false state, which makes the observed actions more informative in expectation. Nonetheless, Theorem 4.14 confirms that $\Theta(\sqrt{n})$ is the fastest achievable convergence rate under any distribution of privacy budgets, and thus represents the optimal learning rate under privacy constraints in sequential settings.